# Scalable Dynamic Origin-Destination Demand Estimation Enhanced by High-Resolution Satellite Imagery Data

### v1

@@ 4.1 Toy network with synthetic data
We begin with a synthetic toy network to evaluate the effectiveness of incorporating density observations into DODE. The network, shown in Figure 4 is modified from Zhang and Qian [2020] which consists of 18 links, including 12 road segments and 6 OD connectors. Detailed link configurations an be found in Table 5. Time-varying traffic conditions, including traffic counts, travel times and densities, are generated based on synthetic demand and recorded at 15-minute intervals to serve as ground truth. As the description in Section 3, we aggregate moving and parking vehicles together to calculate the total remaining vehicles around the segment to represent link densities. To simulate measurement noise, traffic count and travel time observations are corrupted by multiplying by random error terms $\epsilon_{10\%}$ drawn from uniform distribution $\text{Unif}(0.9,1.1)$. The densities are also corrupted by random multiplicative errors of three levels with corresponding uniform distribution: (1) ±$\pm 10\%$ with $\epsilon_{10\%}\sim\text{Unif}(0.9,1.1)$ (2) ±$\pm 20\%$ with $\epsilon_{20\%}\sim\text{Unif}(0.8,1.2)$ and (3) ±$\pm 50\%$ with $\epsilon_{50\%}\sim\text{Unif}(0.5,1.5)$.
To simulate the sparsity of real-world observations, 6 out of 12 links are randomly selected to form the observed link set, denoted by $A_{obs}$. The remaining 6 links are out-of-sample (OOS) links, denoted by link set $A_{oos}$. To prove incorporating density snapshots of the whole network can enhance overall DODE performance, we conduct two separate DODE tests under two data scenarios and compare estimation accuracy on both $A_{obs}$ and $A_{oos}$. In scenario 1, we use only traffic counts and travel times from $A_{obs}$, represented as $\{x_{a}^{t},h_{a}^{t}\}_{t\in T_{d},a\in A_{obs}}$ to mimic the DODE process using localized sensor data. In scenario 2, the data is augmented with 10 density snapshots of all 12 links, which are taken at 15-minute intervals (same as traffic count and travel time data), denoted by $\{k_{a}^{t}\}_{t\in T_{d},a\in A}$. We use coefficient of determination, denoted by $R^{2}$, to measure estimation accuracy of traffic conditions throughout this paper. The main difference of this test between Liu et al. [2024] is incorporating parking information in the link density observations.
Figure 5, 6, 7 and 8 show the count and travel time estimation for unobserved and observed links under different scenario settings. It can be seen that augmenting data with traffic density information for all links can improve the estimation accuracy of both unobserved and observed links, therefore improving the overall performance. However, the improvement relies on the sensing error level when extracting density information from imagery data. The improvement for OOS links are overall better than observed links since density is the only data source for their estimation. Even though the test is performed on a toy network, the DODE is still under-deterministic because only half of links has traffic count and speed observations and the density information cannot differentiate parking and moving vehicles, and these information cannot fully reproduce the truth OD matrix. We calculate the mean squared errors for demand estimation under different scenario settings (shown in Table 2), demonstrating incorporating density can lead to estimated demand which is more closer to ground truth and the improvement of car demand estimation is better than truck due to the relatively higher car volumes on network.
@@
(b) Scenario 2 (±$\pm 10\%$ error in densities)
(c) Scenario 2 (±$\pm 20\%$ error in densities)
(d) Scenario 2 (±$\pm 50\%$ error in densities)
Figure 5: Comparison of count estimation for out-of-sample links
@@
(b) Scenario 2 (±$\pm 10\%$ error in densities)
(c) Scenario 2 (±$\pm 20\%$ error in densities)
(d) Scenario 2 (±$\pm 50\%$ error in densities)
Figure 6: Comparison of travel time estimation for out-of-sample links
@@
Vehicle class Scenario 1 Scenario 2 (±$\pm 10\%$) Scenario 2 (±$\pm 20\%$) Scenario 2 (±$\pm 50\%$)
@@ 4.2.1 Experiment Setting
The numerical experiment uses real-world and synthetic data on Pittsburgh Downtown network to evaluate the proposed framework. The network and sensing scales are shown in Figure 10 and detailed configuration is presented in Table 6. Two distinct real-world data sources are used: traffic count collected by detectors and manual counting provided by the Southwestern Pennsylvania Commission, and two high-resolution satellite images provided by XX. The proposed computer vision pipeline is applied to extract link densities from the satellite images, distinguishing between vehicle classes but not parking and through vehicles. Therefore the resultant link density observations represent total vehicles of both parking and through traffic.
@@ 4.2.2 Traffic density observations from satellite imagery
Two 30-cm resolution satellite images captured by a Pleiades Neo satellite from Airbus were used as the input for the computer vision pipeline. The images were collected on Wednesday, May 11th, 2022 at 12:18 PM and Monday, March 6th, 2023 at 12:17 PM and in Eastern Time. As shown in the first column of images in Figure 12, each image has a width of 5km and a height of approximately 4 kilometers, spanning around 20 square kilometers of surface and covering most of the downtown area in Pittsburgh, PA. The original images included four channels, but only the RGB channels were extracted to run the computer vision pipeline.
To improve object detection results with Airbus imagery, the object detector is also trained with 30-cm resolution satellite imagery from the xView dataset [Lam et al., 2018]. Using satellite imagery with the same resolution for training and inference has been shown to improve object detection results [Van Etten, 2018]. To ensure adequate representation of the positive classes, only the top 10 percent of images with the highest number of annotated vehicles were selected. This subset was then divided into training and validation sets, consisting of 46 images (80%) and 13 images (20%), respectively.
@@
Figure 11 shows vehicle detection, classification and matching results for image patches with a height and width of 100 meters. To analyze results across different types of streets, image patches were collected in areas containing local roads and highways. The yellow bounding boxes in the second column show the vehicle detection output. The third column shows the classification results, with purple and blue boxes representing cars and trucks, respectively. In the fourth column, the red points are the vehicles matched to the street network, while the yellow points represent vehicles that were detected but not matched to the street networks. To exclude vehicles far from the road segments in the network, only vehicles near the road centerline are matched to the street network. Therefore, as shown in the image of the local road in the bottom right corner, vehicles parked near buildings or houses are not matched to the street network.
@@
Figure 12 shows the traffic density at each link of the street network disaggregated by trucks and cars, with the left column showing the satellite images collected for each year. Links with the darkest red color represent traffic densities higher than 40 vehicles per lane-kilometer and are associated with high congestion levels. The traffic density patterns for cars and trucks are consistent across images, which is expected given that both images were collected during the same hour of the day and on weekdays. In addition, the estimated traffic density is higher for cars than for trucks, which is in line with the share of cars and trucks observed in real-world transportation networks.
@@
After computer vision processing, two class-specific density snapshots are obtained within the same time interval (12:15-12:30 PM) on different dates. Figure 23 illustrates the differences between the two data samples. To mitigate the inconsistency, a threshold-based approach is adopted, where links with a difference of at most 5 are selected as consistent links, and their densities are used in the experiments. As a result, 900 out of 1044 links are considered to be consistent, shown as red dots in Figure 23.
@@
To address the inconsistency introduced by multi-source data (e.g., different data providers and different sensing errors in data collection), we propose a two-stage calibration process using two satellite images separately. The first stage calibration is performed using one density snapshot and traffic count data, and we select a subset of links based on the following criteria: for links with non-zero density observations, we selected the link with estimated density no more than three times the observed value for both vehicle classes; for links with zero density observations, we select link with estimated densities no more than 10. As a result, 457 out of 900 link are selected for the second stage. In the second stage, we use the second densities of the selected link subset to perform a separate calibration together with the same traffic count data. This two-stage calibration process helps mitigate the consistency between data sources and leads to more reliable results.
4.2.3 Results and discussion
@@
To evaluate performance of adding additional density data in DODE, we design two DODE scenarios similar to toy network, but this time we only have information of observed links and the ground truth of unobserved links remains hidden. Scenario 1 uses only localized traffic count and speed data of observed links and scenario 2 uses both local sensors and network-wide density snapshot. As DODE is a widely acknowledged underestimate problem with unknown true OD demand and traffic states of unobserved links, we propose the following conditions to be satisfied, justifying the claim that incorporating satellite imagery can enhance DODE performance: (1) adding density into DODE cannot do harm to other data sources, meaning that the accuracy of other traffic condition estimation cannot be largely reduced. (2) the accuracy of density estimation should be improved for observed links, i.e., larger R-squared values for link with density observations. (3) the estimated OD demand pattern in scenario 2 is different from the one in scenario 1.
@@
Figure 13 shows the convergence curves for both DODE tests of two scenarios. Both curves converge after around 70 epochs and the total reduction is similar, indicating that adding additional density loss term in the objective function of DODE will not do harm to the convergence performance. We further plot the estimation results of link-level conditions. Figure 14 shows the accuracy of estimated conditions in DODE without density. It can be seen that link flow estimations match observations well with R-squared values of above 0.9 but using the same network OD flow pattern cannot reproduce the densities. However, when we incorporate density observations in the DODE, the results (Figure 15) show that the estimated flows still match observed traffic counts (with a slight reduction in accuracy) and more importantly, the estimated density can match with observations with much higher accuracy. This can be explained by the under-deterministic nature of DODE problem. Utilizing partial observations (i.e., local sensor data) cannot guarantee the uniqueness of OD matrix. The different demand patterns might lead to the same local traffic conditions, resulting in highly biased estimation, which is difficult to recognize. Adding more data source (i.e., density) is adding more constraints in the problem, leading to smaller feasible region and the estimation results can be closer to the true values. We plot the estimated OD patterns of these two DODE runs and observe some discrepancies for some OD pairs in Figure 16 and 17. The estimation results with density observation satisfy the proposed conditions and we can say incorporating additional density observations can enhance DODE performance.
@@
We further investigate how adding density observations will impact the accuracy of the same localized traffic flow estimation due to the ”trade-off” of fitting multi-source data. We focus on ten overlapping links with both traffic count and density observations and plot the estimation results in two scenarios. Because the satellite image only provide one density snapshot which is in the time interval of 12:15-12:30 PM, we compare the traffic flow estimation in this interval, shown in Figures 18 and 19. It shows that the link flow estimations do not have large discrepancies between two scenarios while truck flow estimations performs worse than cars especially for observed flows that are smaller than 10. This is due to the stochasticity in DNL model and relatively low detection accuracy for trucks in CV pipeline. The unknown inconsistencies between data sources also contribute to this difference.

### v2

@@ 4.1 Toy network with synthetic data
We begin with a synthetic toy network to evaluate the effectiveness of incorporating density observations into DODE. The network, shown in Figure 4 is modified from Zhang and Qian [2020] and consists of 18 links, including 12 road segments and 6 OD connectors. Detailed link configurations can be found in Table 5. Time-varying traffic conditions, including traffic counts, travel times and densities, are generated based on synthetic demand and recorded at 15-minute intervals to serve as ground truth observations. As described in Section 3, we aggregate moving and parking vehicles together to calculate the total remaining vehicles on the segment to represent link densities.
To simulate measurement noise, traffic count and travel time observations are corrupted by multiplying by random error terms $\epsilon_{10\%}$ drawn from a uniform distribution $\text{Unif}(0.9,1.1)$. The densities are also corrupted by random multiplicative errors of three levels with corresponding uniform distributions: (1) $\pm 10\%$ with $\epsilon_{10\%}\sim\text{Unif}(0.9,1.1)$ (2) $\pm 20\%$ with $\epsilon_{20\%}\sim\text{Unif}(0.8,1.2)$ and (3) $\pm 50\%$ with $\epsilon_{50\%}\sim\text{Unif}(0.5,1.5)$.
To simulate the sparsity of real-world observations, 6 out of 12 links are randomly selected to form the observed link set, denoted by $A_{obs}$. The remaining 6 links are unobserved by local sensors, denoted by the link set $A_{unobs}$. To show that incorporating density snapshots of the whole network can enhance overall DODE performance, we conduct two separate DODE tests under two data scenarios and compare estimation accuracy on both $A_{obs}$ and $A_{unobs}$. In scenario 1, we use only traffic counts and travel times from $A_{obs}$, represented as $\{x_{a}^{t},h_{a}^{t}\}_{t\in T_{d},a\in A_{obs}}$ to mimic the DODE process using localized sensor data. In scenario 2, the data is augmented with 10 density snapshots of all 12 links, taken at 15-minute intervals (the same as those of traffic count and travel time data), denoted by $\{k_{a}^{t}\}_{t\in T_{d},a\in A}$.
@@
(b) Scenario 2 ($\pm 10\%$ error in densities)
(c) Scenario 2 ($\pm 20\%$ error in densities)
(d) Scenario 2 ($\pm 50\%$ error in densities)
Figure 5: Comparison of count estimation for unobserved links
@@
(b) Scenario 2 ($\pm 10\%$ error in densities)
(c) Scenario 2 ($\pm 20\%$ error in densities)
(d) Scenario 2 ($\pm 50\%$ error in densities)
Figure 6: Comparison of travel time estimation for unobserved links
@@
Vehicle class Scenario 1 Scenario 2 ($\pm 10\%$) Scenario 2 ($\pm 20\%$) Scenario 2 ($\pm 50\%$)
@@
The full OD ground truth is usually not available, which is typical in real-world networks. Therefore, in this study, the calibration and validation of the estimated multi-class OD demand in all experiments consistently rely on all observed traffic states (which can be either real-world data or synthetic data) that can be reproduced by the network model with estimated OD demand as inputs. Specifically, for each vehicle class, we compare modeled states (e.g., link flow, travel time, density) against observations and compute $R^{2}$ metrics to quantify the goodness-of-fit.
Figure 5, 6, 7 and 8 show the count and travel time estimation for unobserved and observed links under different scenario settings. It can be seen that augmenting the data with traffic density information for all links can improve the estimation accuracy of both unobserved and observed links, therefore improving the overall performance. However, the improvement relies on the sensing error level when extracting density information from imagery data. The improvement for unobserved links is overall better than that for observed links since density is the only data source for their estimation. Even though the test is performed on a toy network, the DODE is still underdetermined because only half of the links have traffic count and speed observations and the density information cannot differentiate parking and moving vehicles, and this information cannot fully reproduce the ground truth OD matrix. We calculate the mean squared errors for demand estimation under different scenario settings (shown in Table 2), demonstrating incorporating density can lead to estimated demand that is closer to ground truth, and that the improvement of car demand estimation is better than that for trucks due to the relatively higher car volumes in the network.
@@ 4.2.1 Experiment Setting
The numerical experiment uses real-world and synthetic data on the Pittsburgh network to evaluate the proposed model. The network and sensing scales are shown in Figure 10 and the detailed configuration is presented in Table 6. Two distinct real-world data sources are used: detector-based traffic counts and manual counts provided by the Southwestern Pennsylvania Commission, and two high-resolution satellite images provided by Airbus. The proposed computer vision pipeline is applied to extract link densities from the satellite images, distinguishing between vehicle classes but not parking and on-road vehicles. Therefore, the link density observations represent total vehicles of both parking and through traffic.
Due to data availability, two satellite images months apart are used in this real-world experiment. Our experiment therefore serves as a proof-of-concept demonstration under limited data availability, to show the benefit of incorporating satellite-derived, network-wide, class-specific density information into the DODE process. When more frequent and temporally aligned imagery data become available, more experiments can be conducted to support large-scale deployment in real world applications.
@@ 4.2.2 Traffic density observations from satellite imagery
Two 30-cm resolution satellite images captured by a Pleiades Neo satellite from Airbus were used as the input for the computer vision pipeline. The images were collected on Wednesday, May 11, 2022 at 12:18 PM and Monday, March 6, 2023 at 12:17 PM Eastern Time. As shown in the first column of images in Figure 12, each image has a width of 5 kilometers and a height of approximately 4 kilometers, spanning around 20 square kilometers and covering the downtown area in Pittsburgh, PA. The original images included four channels, but only the RGB channels were extracted to run the computer vision pipeline.
To improve object detection results with Airbus imagery, the object detector was also trained with 30-cm resolution satellite imagery from the xView dataset [Lam et al., 2018]. Using satellite imagery with the same resolution for training and inference has been shown to improve object detection results [Van Etten, 2018]. To ensure adequate representation of the positive classes, only the top 10 percent of images with the highest number of annotated vehicles were selected. This subset was then divided into training and validation sets, consisting of 46 images (80%) and 13 images (20%), respectively.
@@
Figure 11 shows vehicle detection, classification, and matching results for image patches with a height and width of 100 meters. To analyze results across different types of streets, image patches were collected in areas containing local roads and highways. The yellow bounding boxes in the second column show the vehicle detection output. The third column shows the classification results, with purple and blue boxes representing cars and trucks, respectively. In the fourth column, the red points are the vehicles matched to the street network, while the yellow points represent vehicles that were detected but not matched to the street network. To exclude vehicles far from the road segments in the network, only vehicles near the road centerline are matched to the street network. Therefore, as shown in the image of the local road in the bottom right corner, vehicles parked near buildings or houses are not matched to the street network.
@@
Figure 12 shows the traffic density at each link of the street network, disaggregated by trucks and cars, with the left column showing the satellite images collected for each year. Links with the darkest red color represent traffic densities higher than 40 vehicles per lane-kilometer and are associated with high congestion levels. The traffic density patterns for cars and trucks are consistent across images, which is expected given that both images were collected during the same hour of the day and on weekdays. In addition, the estimated traffic density is higher for cars than for trucks, which is in line with the share of cars and trucks observed in real-world transportation networks.
@@
After computer vision processing, two class-specific density snapshots are obtained within the same time interval (12:15-12:30 PM) on different dates. Figure 26 illustrates the differences between the two samples. To mitigate the inconsistency, a threshold-based approach is adopted, where links with a difference of at most 5 are selected as consistent links, and their densities are used in the experiments. As a result, 900 out of 1044 links are considered consistent, shown as red dots in Figure 26.
@@
To address the inconsistency introduced by multi-source data (e.g., different data providers and different sensing errors in data collection), we propose a two-stage calibration process using two satellite images separately. The first stage calibration is performed using one density snapshot and traffic count data, and we select a subset of links based on the following criteria: for links with non-zero density observations, we select links with estimated density no more than three times the observed value for both vehicle classes; for links with zero density observations, we select links with estimated densities no more than 10. As a result, 457 out of 900 links are selected for the second stage. In the second stage, we use the second densities of the selected link subset to perform a separate calibration together with the same traffic count data. This two-stage calibration process helps mitigate the inconsistencies between data sources and leads to more reliable results.
In practice, satellite imagery-based detection may be affected by different scenes such as network topology, traffic conditions, weather, lighting, occlusions etc. In future study, the CV pipeline can be further enhanced considering these complex conditions, and alternative or improved CV models can be substituted directly without changing the DODE formulation.
4.2.3 Results and Discussions
@@
To evaluate the performance of adding additional density data in DODE, we design two DODE scenarios similar to the toy network, but this time we only have information of observed links and the ground truth of unobserved links remains hidden. Scenario 1 uses only localized traffic counts for observed links, and scenario 2 uses both local sensors and a network-wide density snapshot. As DODE is a widely acknowledged underdetermined problem with unknown true OD demand and traffic states of unobserved links, we propose the following conditions to be satisfied, justifying the claim that incorporating satellite imagery can enhance DODE performance: (1) adding density into DODE cannot do harm to other data sources, meaning that the accuracy of other traffic condition estimation cannot be largely reduced. (2) the accuracy of density estimation should be improved for observed links, i.e., larger R-squared values for links with density observations. (3) the estimated OD demand pattern in scenario 2 is different from the one in scenario 1.
@@
Figure 13 shows the convergence curves for both DODE tests in the two scenarios. Both curves converge after around 70 epochs and the total reduction is similar, indicating that adding an additional density loss term in the objective function of DODE will not hurt the convergence performance. We further plot the estimation results of link-level conditions. Figure 14 shows the accuracy of estimated conditions in DODE without density. It can be seen that link flow estimations match observations well, with R-squared values of above 0.9, but using the same network OD flow pattern cannot reproduce the densities.
However, when we incorporate density observations in the DODE, the results (Figure 15) show that the estimated flows still match observed traffic counts (with a slight reduction in accuracy) and more importantly, the estimated density can match observations with much higher accuracy. This can be explained by the underdetermined nature of DODE problems. Utilizing partial observations (i.e., local sensor data) cannot guarantee the uniqueness of the OD matrix. The different demand patterns might lead to the same local traffic conditions, resulting in highly biased estimation, which is difficult to recognize. Adding a new data source (i.e., density) adds more constraints to the problem, leading to a smaller feasible region and estimation results can be closer to the true values. We further plot the estimated OD patterns of these two DODE runs and observe some discrepancies for some OD pairs in Figure 16 and 17. This experiment serves as a self-benchmarking ablation analysis to isolate the marginal contribution of satellite-derived density. The increased accuracy of density estimation demonstrates the value added by satellite imagery as a complementary data source.
@@
We also investigate how adding density observations impacts the accuracy of the same localized traffic flow estimation due to the ”trade-off” of fitting multi-source data. We focus on ten overlapping links with both traffic count and density observations and plot the estimation results in two scenarios. Because the satellite image only provides one density snapshot in the time interval of 12:15-12:30 PM, we compare the traffic flow estimation in this interval, shown in Figures 18 and 19. The results show that the link flow estimations do not have large discrepancies between the two scenarios, while truck flow estimations perform worse than cars especially for observed flows that are smaller than 10. This is due to the stochasticity in the DNL model and the relatively low detection accuracy for trucks in the CV pipeline. Unknown inconsistencies between data sources also contribute to this difference.
From a computational cost perspective, incorporating density data does not change the main structure of the CG-based DODE framework. The dominant computation cost still comes from DNL runs in the forward pass that generates traffic state variables, and the computation of gradient backpropagation in the backward pass. Density is one of the state variables produced by the DNL model (extracted from cumulative curves), and its gradient is computed using the same set of DAR matrices that map path flows to link inflow/outflow. Therefore, adding the density loss does not introduce much additional computation cost in solving DODE.
Moreover, satellite image processing is performed offline as a preprocessing step that converts raw images into link-level density observations, which can be used in the offline DODE. In the Pittsburgh case, the vehicle detection for each image took approximately 7 minutes end-to-end. In our implementation, the CV preprocessing time is smaller than the time for running DNL per iteration, so the overall runtime is dominated by OD estimation rather than image processing. The CV pipeline uses a sliding-window (tiling) strategy, so its complexity can be characterized by the number of tiles needed to cover the spatial extent of the network. The total inference cost in the CV pipeline grows roughly linearly with the area covered by the satellite imagery. Importantly, scaling to larger networks does not increase per-tile complexity and instead only increases the number of tiles. This makes the preprocessing step easy to parallelize across CPU cores or GPUs, and straightforward to scale using batch inference or distributed execution when needed.
@@
4.2.4 Comparison with the low rank-based method
We use the Principal Component-based Simultaneous Perturbation Stochastic Approximation (PC-SPSA) method proposed by Qurashi et al. [2019, 2022] as a benchmark. PC-SPSA first learns a low-dimensional representation of dynamic OD demand by extracting principal components (PCs) from a set of historical OD samples, and then performs OD estimation in this reduced PC space. This benchmark represents a different line of work from our approach, as described in the literature review section. Our CG-based framework mitigates the underdetermined issue by incorporating heterogeneous, independent observations (traffic counts and satellite-derived densities) as regularization, whereas PC-SPSA relies on a low-rank assumption to reduce the DODE problem dimension via PCA.
In the benchmark experiment, we use the same real-world observations (traffic counts and satellite-derived densities) as in our proposed model for PC-SPSA. We also follow the same initialization process by starting from a randomly generated dynamic OD matrix rather than a pre-specified seed OD. Because the two methods do not necessarily start from the same random point, their absolute objective values are not directly comparable across runs. Therefore, we report the relative objective reduction (normalized by the initial objective value) to show the convergence behaviors.
For the Pittsburgh network, historical dynamic OD matrices are not available, so the low-rank structure of dynamic OD demand required by PCA cannot be directly inferred from real observations. To construct a feasible PCA-based benchmark, we generate a set of OD samples following the synthetic data generation procedure in Qurashi et al. [2022]. Specifically, we first obtain one dynamic OD matrix from a prior DODE run and use it as a template to generate additional OD samples (i.e., it is not used as the initialization in solving DODE with PC-SPSA). We generate 299 additional dynamic OD matrices, each covering a 2-hour time horizon (8 intervals at 15-minute resolution). This yields 300 OD matrices in total, corresponding to 2400 interval-level OD vectors for each vehicle class, which are used to compute the PCs. We compare the two methods in terms of (1) convergence speed, assessed by the normalized objective function reduction over the same number of iterations, and (2) goodness-of-fit between estimated and observed traffic states, quantified by $R^{2}$ for class-specific link flows and densities.
Results show that PC-SPSA underperforms our method on the Pittsburgh network in both convergence speed (Figure 20) and state estimation accuracy (Figures 21 and 22). This finding is consistent with Ma et al. [2020] who reported worse performance of the SPSA method compared with CG-based approaches on a large-scale network. We explain this performance gap from three perspectives. First, the SPSA method estimates gradients of the objective function via stochastic perturbations, while our method computes exact gradients with respect to flow and density through the DAR matrices after each DNL run, yielding a deterministic and interpretable mapping from OD flows to link-level traffic states. Although PCA reduces the DODE problem to a lower dimension, the stochastic-gradient nature of SPSA can still slow convergence and degrade fitting, especially for density observations that couple in and out link flows across multiple time intervals through accumulation and propagation effects. Second, without real-world historical OD data, it remains unclear whether the true OD patterns in Pittsburgh have a low-rank structure. If this assumption is not valid, PCA-based compression may discard spatio-temporal variations of OD demand and degrade the estimation. Third, PC-SPSA performance is sensitive to hyperparameter choices; despite testing settings reported in Qurashi et al. [2019, 2022] and additional tuning on our dataset, improvements were limited and more extensive tuning may be required to yield better performance. Overall, our method yields better performance on the Pittsburgh network and we leave more extensive benchmarking experiments to future work when richer real-world historical OD datasets become available.
Figure 20: Normalized objective reduction (relative to the initial objective) for CG-based method vs. PC-SPSA on the Pittsburgh network
Figure 21: Comparison for class-specific link flow estimation (CG-based vs. PC-SPSA)
Figure 22: Comparison for class-specific link density estimation (CG-based vs. PC-SPSA)